OpenStudy (anonymous):

If g(x)=-x^2-10 find g(-4)

OpenStudy (anonymous):

well g(-4)= -(-4)^2-10 = -16-10=-26

OpenStudy (anonymous):

g(-4) means to find the value of the function -x^2 - 10 when x = -4